### v1

@@ Results
MIRO offers the capability of simultaneously handling diverse structural patterns, by compressing localizations from different cluster shapes into a uniform representation. This capability enables effective clustering using a single set of parameters across different shapes when applied to algorithms like DBSCAN. The unified representation simplifies parameter tuning and enhances clustering performance. However, this transformation can also lead to challenges in subsequent classification, as the uniform collapse of different shapes may obscure their unique features.
@@
However, while transforming various structures into compact forms, MIRO generates additional information that can be encoded as node features to enable simultaneous shape classification. This dual capability is essential for, e.g., distinguishing among various molecular assemblies within the same biological environment, each exhibiting unique organizational patterns and functional roles.
@@
To evaluate MIRO’s ability to simultaneously cluster and classify different structures, we generated simulated datasets comprising mixtures of circular, elliptical, C-shaped, and ring-shaped clusters (Figure 3). Each cluster type represents a distinct molecular assembly, characterized by unique spatial properties. MIRO effectively captures these spatial features, enabling accurate separation and identification of different structures within a single dataset.
@@
Figure 3 illustrates the results for three distinct mixtures: spots and ellipses (Figures 3a–c), spots and rings (Figures 3d–f), and C-shaped clusters and rings (Figures 3g–i). Overall, the results demonstrate that MIRO’s preprocessing effectively distinguishes between different shapes and accurately assigns localizations to their respective clusters. This enhanced performance is evident in both the confusion matrices (Figures 3b, 3e, and 3h), which show higher classification accuracy across all shape combinations, and the clustering metrics (Figures 3c, 3f, and 3i). Notably, the clustering metrics indicate that, in several instances, the performance is similar to those obtained for a single shape. This is particularly remarkable considering that no restrictions were imposed on cluster overlap; clusters of different shapes could overlap or be arranged in ways that mimic other shapes, such as aligned spots forming an ellipse or facing C-shapes resembling a ring.
Detecting heterogeneous and dense clusters
@@
MIRO offers substantial improvements for analyzing adjacent clusters in SMLM data. We assessed MIRO’s effectiveness by conducting quantitative tests as a function of the inter-cluster distance. We simulated pairs of clusters with similar sizes but containing different numbers of localizations, located at varying cluster-to-cluster distances. Localizations belonging to the same cluster were spatially arranged according to a 2D Gaussian distribution with width . The number of localizations per cluster was drawn from an exponential distribution. Clusters were spaced at various distances as a function of . We applied MIRO and DBSCAN to compare the methods’ ability to resolve the clusters, as quantified by the Jaccard Index for cluster detection (JI). As demonstrated in Figure 4a, at distances , MIRO significantly improves clustering accuracy compared to DBSCAN, providing a more precise characterization of nanocluster spatial arrangements and thus improving their quantification.
@@
Multiscale clustering of nuclear pore complex
@@
To demonstrate MIRO’s ability to tackle these challenges simultaneously and quantitatively, we first relied on simulations. We generated synthetic localization maps with structures composed of small symmetrical clusters, each with a random number of localizations, arranged in rings with eightfold symmetry. As shown in Figure 1, the MIRO architecture was trained to collapse localizations forming the spots and the ring-shaped clusters toward their respective centers. The results of the ensuing clustering, shown in Figures 5a-b, demonstrate that MIRO can work simultaneously across multiple scales. Specifically, MIRO provides significant performance enhancements compared to DBSCAN at both the ring and spot scales, achieving better scores in all metrics.
@@
The quantitative analysis of the clustering results is presented in Figures 5e-f. The histogram of the number of localizations per NPC in Figure 5e shows very good agreement with the one obtained in the original article (Figure 4g in Ref. [13]), where segmentation was performed using a specifically designed algorithm involving multiple filtering and thresholding of reconstructed super-resolution images. Similarly, Figure 5f presents a histogram of the number of localizations per spot, revealing an exponential distribution with an average of 12.28 localizations per spot. Considering that each corner hosts 4 Nup96 proteins, this result corresponds to approximately 3 localizations per protein, closely aligning with the estimation performed in the original article (2.8 localizations, Table 1 in Ref. [13]). These analyses thus confirm the effectiveness of MIRO compared to the ad hoc analysis workflow used in the original study.
Discussion
@@
Preprocessing SMLM datasets with MIRO enhances the performance of algorithms for complete clustering. Accurate clustering enables the quantitative assessments of spatial organization within cellular environments, through the precise estimation of quantities such as cluster size, protein copy number, and intercluster distances, leading to deeper insights into biological phenomena [14, 12, 15].
The integration of MIRO allows for simultaneous clustering and classification of various structural patterns within a single dataset. To the best of our knowledge, this feature is not offered by any of the previous methods. Moreover, MIRO operates in both multiclass and multiscale modalities, with the multiscale approach being particularly advantageous for nested structures, such as NPCs [13]. It is important to note that the same MIRO block is used to compress structures of different sizes and shapes in a single forward pass, thus the latent representation inherently learns hierarchies and scales within the data.
MIRO advances the field of data-driven approaches for SMLM by offering augmented functionality. Earlier data-driven methods, such as those based on recurrent neural networks, were limited to binary classification, distinguishing only between clustered and non-clustered localizations [20]. More recently, graph neural networks have shown promise in clustering SMLM data with simple symmetric shapes, but they do not perform any further classification [22]. Additionally, unsupervised methods like SEMORE [40] rely on DBSCAN for clustering, which could be improved by incorporating MIRO, and only use machine learning for cluster classification.
@@
From an architectural point of view, MIRO tackles several technical challenges and introduces an innovative scheme for the application of rGNNs in the analysis of point clouds. A primary challenge addressed by MIRO is handling the high density of localizations, which requires a wide receptive field to capture spatial relationships within dense point clouds [41]. This can be computationally expensive and resource-intensive. MIRO addresses this by employing a recurrent structure that amplifies the receptive field while maintaining a limited number of connections, significantly reducing computational costs and training parameters.
A key feature of MIRO’s architecture is its approach to message passing [29]. Unlike standard techniques where nodes retain and aggregate previous features at each step, MIRO’s hidden nodes mainly focus on spatial relationships. This design prevents the over-smoothing of node features [41]. In addition, the input graph does not update the semantic information and preserves topological information crucial for cluster recognition at every step.
While MIRO offers significant advantages, it is not without limitations. One fundamental challenge of MIRO, and all clustering methods, is accurately identifying and separating structures overlapping with either noise or other structures. Future improvements in this sense will be crucial for advancing clustering methods in complex biological datasets. Help in this sense might come from extending the node features. Node features in MIRO can encompass a wide range of attributes, providing flexibility in data representation. While our current implementation does not utilize temporal information, incorporating such data in an embedded form could enrich the model’s performance by accounting for photophysical effects [42, 43].
Online Methods
MIRO’s graph representation
The input to MIRO is a graph representation [33] of an SMLM point cloud (Figure 1a–c). In this graph, nodes () represent individual molecular localizations and edges () capture the spatial relationships within the point cloud derived from a Delaunay triangulation. To ensure that only meaningful, local spatial relationships are retained, the edges are filtered based on a distance threshold selected according to the local density of the point cloud.
@@
Edge features encode relational attributes between nodes and , such as the Euclidean distance and positional displacement describing their relative arrangement. This selection of node and edge features allows MIRO to inherently analyze graphs of varying sizes and spatial extents without introducing additional processing complexities. Nevertheless, there are no intrinsic restrictions on the type or number of descriptors that can be encoded in the node and edge features.
MIRO’s architecture
@@
MIRO transforms the input node and edge features into higher-level latent representations through learnable linear transformations, mapping and into latent vectors and , each with a dimensionality of 256. This latent representation serves as the input to a sequence of blocks that recurrently update a hidden graph . At each step, the updated hidden nodes features are also decoded through learnable linear transformations and used to calculate the loss function (Figure 1c).
@@
The core operations are executed within the MIRO block (Figure 1d). At each step , concatenates the latent representations and with the node and edges of the hidden graph , producing and (Figure 1d). The hidden graph features and are initialized as zeros and, as the recursive process unfolds, they are progressively refined as
@@
where denotes concatenation, is the neighborhood of node , and the functions and represent linear layers followed by a ReLU activation, which map the output into a 256-dimensional space.
@@
In these operations, the hidden representations play a crucial role in progressively refining the understanding of each node’s context within the graph. It is important to note that, for updating the hidden node features , we purely rely on the updated edge hidden states , without including skip connections to the current node hidden states, which is common in various flavors of message passing [29]. The rationale for this choice is to better equip MIRO to discern and emphasize the immediate structural context of each node. This choice has resulted in a substantial improvement in training stability.
@@
At each step, the MIRO block further uses a learnable linear transformation to decode the updated hidden node features , generating a displacement vector in Cartesian space for each molecular localization. The objective of these learned displacements is, when summed with the localization coordinates, to shift localization belonging to the same cluster toward a common center, resulting in a compact representation of clusters within the SMLM point cloud, while leaving background localizations unaltered.
Training MIRO
@@
To ensure that the hidden graph representation remains meaningful and to prevent vanishing gradients, the loss is calculated at each recurrent step of the model. This approach implicitly imposes regularization on the displacement vectors, helping to maintain the structural integrity of the clusters throughout the training process. By calculating the loss at multiple stages, MIRO is encouraged to refine the displacement vectors incrementally, preventing early steps from degrading the quality of later predictions and ensuring consistent optimization across the entire sequence of recurrent updates.
@@
where represents the set of all pairs of neighboring nodes, denotes the Euclidean distance function, and and describe the shifted positions after adding the predicted and target displacement to the original localizations. Although both loss functions aim to achieve a similar outcome, we observe that their combined application enhances the model’s ability to form compact and well-defined clusters.
@@
where is the predicted class label for node at recurrent step , represents the true class label for node , and is a weighting factor.
Dataset Augmentation
MIRO can be effectively trained using a small set of representative clusters , typically on the order of a few tens. As shown in the example provided at https://github.com/DeepTrackAI/MIRO/, reasonable results can be obtained with .
To learn to generalize over different cluster arrangements, the framework uses the clusters to generate an augmented dataset of point clouds by applying a series of transformations. Augmented clusters are generated by randomly applying rotations, reflections, localization dropouts/additions, and localization displacements. These augmented clusters are then randomly placed within a field-of-view to generate training point clouds. Background localizations are further added to each generated point cloud, either from a uniform spatial distribution (non-blinking) or by sampling background localizations from the benchmark study data, augmented as described above for clusters (blinking). The number of training clusters and point clouds necessary to achieve specific performance levels vary with cluster shape and density.

### v2

@@ 3.3 Simultaneous clustering and classification of different shapes
MIRO offers the capability of simultaneously handling diverse structural patterns by compressing localizations from different cluster shapes into a uniform representation. This capability enables effective clustering using a single set of parameters across different shapes when applied to algorithms like DBSCAN. The unified representation simplifies parameter tuning and enhances clustering performance. However, this transformation can also lead to challenges in subsequent classification, as the uniform collapse of different shapes may obscure their unique features.
@@
However, while transforming various structures into compact forms, MIRO can generate additional output features at the node level that can be used, e.g., for simultaneous cluster shape classification. This dual capability is essential for, e.g., distinguishing among various molecular assemblies within the same biological environment, each exhibiting unique organizational patterns and functional roles.
@@
To evaluate MIRO’s ability to simultaneously cluster and classify different structures, we generated simulated datasets comprising mixtures of circular, elliptical, C-shaped, and ring-shaped clusters (Figure 3). Each cluster type represents a distinct molecular assembly, characterized by unique spatial properties. MIRO effectively learns to capture these features at the node level. While clustering ensures accurate separation of structures, taking the mode of node-level class predictions within each cluster allows for reliable identification of the corresponding structural type in heterogeneous datasets.
@@
Figure 3 illustrates the results for three distinct mixtures: spots and ellipses (Figure 3a–c), spots and rings (Figure 3d–f), and C-shaped clusters and rings (Figure 3g–i). Overall, the results demonstrate that MIRO’s preprocessing effectively distinguishes between different shapes and accurately assigns localizations to their respective clusters. This enhanced performance is evident in both the confusion matrices (Figure 3b, e, and h), which show higher classification accuracy across all shape combinations, and the clustering metrics (Figure 3c, f, and i). Notably, the clustering metrics indicate that, in several instances, the performance is similar to those obtained for a single shape. This is particularly remarkable considering that no restrictions were imposed on cluster overlap; clusters of different shapes could overlap or be arranged in ways that mimic other shapes, such as aligned spots forming an ellipse or facing C-shapes resembling a ring.
3.4 Detecting heterogeneous and dense clusters
@@
MIRO offers substantial improvements for analyzing adjacent clusters in SMLM data. We assessed MIRO’s effectiveness by conducting quantitative tests as a function of the inter-cluster distance. We simulated pairs of clusters with similar sizes but containing different numbers of localizations, located at varying cluster-to-cluster distances. Localizations belonging to the same cluster were spatially arranged according to a 2D Gaussian distribution with width . The number of localizations per cluster was drawn from an exponential distribution. Clusters were spaced at various distances as a function of . We applied MIRO and DBSCAN to compare the methods’ ability to resolve the clusters, as quantified by the Jaccard Index for cluster detection (JIc). As demonstrated in Figure 4a, at distances , MIRO significantly improves clustering accuracy compared to DBSCAN, providing a more precise characterization of nanocluster spatial arrangements and thus improving their quantification.
@@
3.5 Multiscale clustering of nuclear pore complex
@@
To demonstrate MIRO’s ability to tackle these challenges simultaneously and quantitatively, we first relied on simulations. We generated synthetic localization maps with structures composed of small symmetrical clusters, each with a random number of localizations, arranged in rings with eightfold symmetry. As shown in Figure 1, the MIRO architecture was trained to collapse localizations forming the spots and the ring-shaped clusters toward their respective centers. The results of the ensuing clustering, shown in Figure 5a–b, demonstrate that MIRO can work simultaneously across multiple scales. Specifically, MIRO provides significant performance enhancements compared to DBSCAN at both the ring and spot scales, achieving better scores in all metrics.
@@
The quantitative analysis of the clustering results is presented in Figure 5e–f. The histogram of the number of localizations per NPC in Figure 5e shows very good agreement with the one obtained in the original article (Figure 4g in Ref. [13]), where segmentation was performed using a specifically designed algorithm involving multiple filtering and thresholding of reconstructed super-resolution images. Similarly, Figure 5f presents a histogram of the number of localizations per spot, revealing an exponential distribution with an average of 12.28 localizations per spot. Considering that each corner hosts 4 Nup96 proteins, this result corresponds to approximately 3 localizations per protein, closely aligning with the estimation performed in the original article (2.8 localizations, Table 1 in Ref. [13]). These results highlight the accuracy and reliability of MIRO in multiscale real-data applications. In contrast, DBSCAN struggles in this scenario. As shown in Supplementary Figure 1, identifying optimal DBSCAN parameters is non-trivial, and the algorithm often produces fragmented clusters, resulting in an artificial peak of small-sized clusters in the size distribution. Overall, these findings underscore the advantage of MIRO not only over traditional clustering approaches like DBSCAN, but also compared to the task-specific, multi-step analysis pipeline originally proposed for this dataset [13].
4 Discussion
@@
Preprocessing SMLM datasets with MIRO enhances the performance of algorithms for complete clustering. Accurate clustering enables the quantitative assessments of spatial organization within cellular environments, through the precise estimation of quantities such as cluster size, protein copy number, and inter-cluster distances, leading to deeper insights into biological phenomena [14, 12, 15].
The integration of MIRO allows for simultaneous clustering and classification of various structural patterns within a single dataset. To the best of our knowledge, this feature is not offered by any of the previous methods. Moreover, MIRO operates in both multiclass and multiscale modalities, with the multiscale approach being particularly advantageous for nested structures, such as NPCs [13]. It is important to note that the same MIRO block is used to compress structures of different sizes and shapes in a single forward pass, thus the hidden representation inherently learns hierarchies and scales within the data.
MIRO advances data-driven analysis in SMLM by offering expanded functionality beyond existing approaches. Early machine learning methods, such as those based on recurrent neural networks, were limited to binary classification, distinguishing only between clustered and non-clustered localizations [20]. More recently, graph neural networks have been used to cluster SMLM data with simple symmetric shapes, but these models do not support classification of structural types [22]. SEMORE [42] introduces a different strategy by applying machine learning for morphological fingerprinting of clusters obtained from density-based clustering. In contrast, MIRO focuses on learning robust spatial representations, enabling both improved clustering and simultaneous structural classification. As such, MIRO could serve as a valuable preprocessing step that complements methods like SEMORE, potentially improving the quality of the point clouds that SEMORE subsequently analyzes.
Although MIRO is trained in a supervised manner, we note that supervision alone does not guarantee improved clustering results — as shown in the benchmark study, where the supervised method CAML [20] does not outperform the best-performing unsupervised alternatives. MIRO’s superior performance stems not simply from supervision, but also from its tailored architecture and inductive biases, which enhance representation quality and generalization even with limited training data.
This point is further reinforced by our comparison with MAGIK-S (see Supplementary Table 6), an implementation of a recent supervised method based on graph neural networks [22]. While both MIRO and MAGIK-S incorporate relational inductive biases through the use of graph-based models, the two architectures are designed to tackle different tasks. MAGIK-S is structured as a two-step pipeline involving node-level and edge-level classification, ultimately relying on community detection to form clusters. This approach requires a more complex optimization process, longer training times, and a greater amount of labeled data. In contrast, MIRO is designed specifically to enhance latent representations through recurrent graph operations, allowing for efficient single-shot or few-shot learning and fast generalization across scenarios.
Moreover, the architectural biases in MIRO are fundamentally different from those in MAGIK-S. MIRO explicitly decouples topological refinement from semantic input by maintaining access to the original graph structure across recurrent steps, which enables it to build robust, scale-adaptive representations. This design leads to improved clustering performance with significantly lower computational and data requirements.
@@
From an architectural point of view, MIRO tackles several technical challenges and introduces an innovative scheme for the application of rGNNs in the analysis of point clouds.
A key challenge addressed by MIRO is the analysis of high-density localization maps, which requires a broad receptive field to capture spatial relationships within dense point clouds. In conventional message-passing architectures, achieving this typically demands deeper networks, resulting in increased computational cost, memory usage, and a higher risk of oversmoothing [43]. MIRO overcomes these limitations by employing a recurrent architecture that progressively expands the receptive field without increasing the number of trainable parameters.
MIRO adopts a non-conventional approach to message passing [30], aiming to progressively refine topological information, captured in the hidden state, while maintaining access to the semantic information encoded in the original graph. To achieve this, MIRO’s update block is specifically designed to focus on the local structural context: hidden node features are updated exclusively from the aggregated messages (i.e., hidden edge features), without concatenating previous node features. Meanwhile, the unmodified input graph is passed through each recurrent step, preserving the original semantic features and ensuring they remain accessible throughout the computation. This architectural choice enables MIRO to disentangle structural and semantic cues, leading to more interpretable representations that are particularly well-suited for identifying spatially coherent clusters.
While MIRO offers significant advantages, it is not without limitations. One fundamental challenge of MIRO, and all clustering methods, is accurately identifying and separating structures overlapping with either noise or other structures. Future improvements in this sense will be crucial for advancing clustering methods in complex biological datasets. Help in this sense might come from extending the node features. Node features in MIRO can encompass a wide range of attributes, providing flexibility in data representation. While our current implementation does not utilize temporal information, incorporating such data in an embedded form could enrich the model’s performance by accounting for photophysical effects [44, 45, 46, 42].
Beyond SMLM, MIRO’s core capabilities make it a promising tool for a range of scientific domains. In neuroscience, for example, MIRO could help delineate complex neural circuits from sparse imaging data, advancing our understanding of brain connectivity [47]. Similarly, in environmental science, it could aid in uncovering spatial patterns in ecological datasets, such as species distributions or pollution gradients [48], contributing to more data-driven environmental monitoring and modeling.
5 Methods
5.1 MIRO graph representation
The input to MIRO is a graph representation [34] of an SMLM point cloud (Figure 1a–c). In this graph, nodes () represent individual molecular localizations and edges () capture the spatial relationships within the point cloud derived from a Delaunay triangulation. To ensure that only meaningful, local spatial relationships are retained, the edges are filtered based on a distance threshold selected according to the local density of the point cloud (Supplementary Table 7).
@@ 5.1 MIRO graph representation
Edge features encode relational attributes between nodes and , such as the Euclidean distance and positional displacement describing their relative arrangement. In the current implementation, each edge feature includes the vector , where and are the coordinates of nodes and , respectively. Although the input graph is undirected — meaning that information flows symmetrically between connected nodes — directional information is preserved by assigning displacement vectors and to each message-passing direction. This formulation enables MIRO to capture relative orientations of neighboring nodes while maintaining the symmetry of the graph structure.
This selection of node and edge features allows MIRO to inherently analyze graphs of varying sizes and spatial extents without requiring additional processing complexity. Moreover, the architecture is agnostic to the specific type or number of descriptors used: while distance and directional cues serve as the primary relational information in our current implementation, the framework can readily incorporate additional edge attributes — such as temporal proximity, semantic labels, topological relations, or domain-specific measurements — depending on the needs of the application.
5.2 MIRO architecture
@@
MIRO transforms the input node and edge features into higher-level latent representations using a learnable dense layer followed by ReLU activation, mapping and into latent vectors and , each with a dimensionality of 256. This latent representation serves as the input to a sequence of recurrent steps that recurrently update a hidden graph . At each recurrent step, the updated hidden nodes features are also decoded through a learnable dense layer and used to calculate the loss function (Figure 1c).
@@
The core operations are executed within the MIRO block (Figure 1d). At each recurrent step , concatenates the latent representations and with the node and edges of the hidden graph , producing and (Figure 1d). The hidden graph features and are initialized as zeros and, as the recursive process unfolds, they are progressively refined as
@@
where denotes concatenation, is the neighborhood of node , and the functions and represent dense layers followed by ReLU activations, which map the output into a 256-dimensional space.
@@
In these operations, the hidden representations play a crucial role in progressively refining the understanding of each node’s context within the graph (see Number of recurrent steps: influence on performance and oversmoothing). It is important to note that, for updating the hidden node features , we purely rely on the updated edge hidden states , without including skip connections to the current node hidden states, which is common in various flavors of message passing [30]. The rationale for this choice is to better equip MIRO to discern and emphasize the structural context of each node.
@@
At each recurrent step, the MIRO block further uses a learnable dense layer to decode the updated hidden node features , generating a displacement vector in Cartesian space for each molecular localization. The objective of these learned displacements is, when summed with the localization coordinates, to shift localization belonging to the same cluster toward a common center, resulting in a compact representation of clusters within the SMLM point cloud, while leaving background localizations unaltered.
5.3 MIRO loss function
@@
To ensure that the hidden graph representation remains meaningful and to prevent vanishing gradients, at each iteration within an epoch, the loss is further averaged across all recurrent steps. This approach implicitly imposes regularization on the displacement vectors, helping to maintain the structural integrity of the clusters throughout the recurrent step. By calculating the loss at all recurrent steps, MIRO is encouraged to refine the displacement vectors incrementally, preventing early steps from degrading the quality of later predictions and ensuring consistent optimization across the entire sequence of recurrent updates.
@@
where is a weighting factor, represents the set of all pairs of neighboring nodes, denotes the Euclidean distance function, and and describe the shifted positions after adding the predicted and target displacement to the original localizations. Although both loss functions aim to achieve a similar outcome, we observe that their combined application enhances the model’s ability to form compact and well-defined clusters.
@@
where is a weighting factor, is the predicted class label for node at recurrent step , and represents the true class label for node .
5.4 MIRO training and augmentations
MIRO supports an effective one- or few-shot learning paradigm, in which training can be performed using as little as a single representative cluster (). This is possible due to two key factors: (i) the weak conservation of cluster shape and spatial organization, and (ii) a systematic dataset augmentation strategy that promotes generalization across spatial contexts and noise conditions. As demonstrated in the results displayed in Supplementary Table 5, MIRO achieves reasonable performance when trained with just one annotated cluster. An end-to-end example of single-shot training is provided in the GitHub repository https://github.com/DeepTrackAI/MIRO/.
To leverage this data efficiency, MIRO employs an augmentation pipeline that transforms the small set of training clusters () into a large number of diverse point clouds () for training. Each point cloud is generated by applying a series of transformations to randomly selected clusters, including geometric transformations (rotations, reflections), stochastic perturbations (localization dropout or addition), and spatial jitter (small random displacements). These transformed clusters are then randomly placed within a synthetic FOV to generate the final training samples.
To further mimic realistic imaging conditions, background localizations are added to each point cloud. For non-blinking scenarios, background points are sampled from a uniform spatial distribution. For blinking scenarios, background is sampled from benchmark data and similarly augmented to reflect variations in localization behavior.
This approach allows the model to learn a robust, transferable representation from minimal annotated input, significantly reducing the need for extensive labeled training data. While the number of clusters and point clouds required depends on shape complexity and density, we found that stable performance could be achieved with small training sets (Supplementary Table 5).